OpenStudy (anonymous):

What is the value of y in the equation 4y – 2(1 – y) = –44?

OpenStudy (anonymous):

4y - 2(1 - y) = -44 4y - 2 + 2y = -44 4y + 2y = -44 + 2 6y = -42 y = -42/6 y = -7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

What is the value of -3|15 - s| + 2s3 when s = -3?